# Tension-induced giant actuation in unstructured elastic sheets

### v1

@@ D.1 The planar state
We benchmark the expressions for the stress in the planar state (A10) with the finite-element simulations (see Appendix B) of a sheet ($L/W=2,~{}t/W=0.0025,~{}W_{\mathrm{load}}/W=0.1,~{}E=3000~{}\text{MPa},~{}\nu=0.3$) subjected to longitudinal strain $u_{y}/L=0.025$. Considering the scaling assumed previously (i.e. lengths are rescaled by $L/2$ and $L=2$, and stresses are normalized by $T=Y\,u_{y}/L$), we plot in Fig. A7 the planar stress state obtained by (A10) and the numerical simulations (for which we consider only the region $x\geq W_{\mathrm{load}}/2$). We notice a good agreement between the analytic calculations (Fig. A7a and d) and the numerics (Fig. A7b and e). The transverse stress (Fig. A7a and b) is tensile at the corners where we exert $T$, and compressive as we move away from the corners towards the centre — we have included the contour line at $\sigma_{xx}/T=0$ to evince the coexistence of compressive and tensile areas in the sheet. At a distance close to the $y$-axis, the cross-sectional profiles of this stress component (Fig. A7c) help us visualize the tension-to-compression transition as we move away from the corners, and the stress singularity at their vicinity (see the next section for more details on the stress singularities at the corners of a sheet). Away from the $y$-axis, the transverse stress within the sheet fades rapidly. Conversely, the longitudinal stress (Fig. A7d and e) is tensile everywhere in the sheet, and hence deserves less attention in our discussion.
@@
Extended Data Fig. A7: Stress in the planar state. a, Analytic solution for the transverse stress (A10a) within the sheet subjected to a longitudinal localized load at the corners as per Fig. A6b; $W/L=0.45$ and we take $n=16$ terms in the series. b, Numeric results for the transverse stress obtained by finite-element simulations (see Appendix B) of a sheet ($L/W=2,~{}t/W=0.0025,~{}W_{\mathrm{load}}/W=0.1,~{}E=3000~{}\text{MPa},~{}\nu=0.3$) subjected to a longitudinal strain $u_{y}/L=0.025$. The numeric results are rescaled according to the scaling in the mathematical model, and they are only shown within $x\geq W_{\mathrm{load}}/2$. The same colorbar as in a applies. c, Cross-sectional profiles of the transverse stress at different values of $x$; symbols for the numerics, lines for the analytic expression. d, Longitudinal component of the stress (A10b). e, Numeric results for the longitudinal stress with the same colorbar as in d.
@@
We further examine the transverse stress for different values of the aspect ratio $W/L$. We already noticed that the transverse stresses are weak away from the corners of the sheet that are loaded. Thus, we focus on the stress state close to the loaded corners. We plot in Fig. A8 the expression (A10a) shifted in $y$, $y\to y+1$, and we scale both axes by $W/2$. We observe a universal stress state for $W/L\lesssim 1/4$: the transverse stress is tensile at the vicinity of the corner, and compressive within an area that borders the narrow tensile part. The width of this compressing sector scales with $\sim W$ — we have included contour lines every $\Delta\sigma_{xx}=-0.05$ in the compressive part so as to evince the behaviour of these stresses. This scaling is paramount in our analysis in the main text. For $W/L\gtrsim 1/4$, the compressive sectors corresponding to both corners overlap, and hence the sheet has a single compressed area spanning the centre of the sheet, while the two tensile sectors prevail at the corners.
@@
Extended Data Fig. A8: Transverse stress in the planar state for different aspect ratios. a-h, Density plots of the stress by the expression (A10a). The longitudinal coordinate, $y$, is shifted $y\to y+1$; and both axes by $W/2$. The same colorbar (only shown in a) applies to all panels. The number of terms, $n$, taken in the series (A10a) is labeled in each graph. Contour lines are displayed at $\sigma_{xx}/T=(0,-0.05,-0.1,-0.15)$.
D.2 Benthem’s Laplace transform method
We define the function $f(x,p)$ as the Laplace transformation of the Airy potential,
$f(y,p)=\int_{0}^{W/2}\chi(x,y)\,e^{-px}\,\mathrm{d}x,$ (A11)
@@ D.2 Benthem’s Laplace transform method
$\chi(x,y)=\frac{1}{2\pi\mathrm{i}}\int_{c-\mathrm{i}\infty}^{c+\mathrm{i}\infty}f(y,p)\,e^{px}\,\mathrm{d}p\quad(c\in\mathbb{R},c>0).$ (A12)
We verify that (A12) is the inverse of (A11). Substituting the integrand in (A12) by the Laplace transform definition (A11) and pursuing some algebraic manipulations, we obtain:
$\chi(x,y)=$ $\frac{1}{2\pi\mathrm{i}}\int_{c-\mathrm{i}\infty}^{c+\mathrm{i}\infty}\mathrm{d}p\,e^{px}\,\int_{0}^{W/2}\chi(\hat{x},y)\,e^{-p\hat{x}}\,\mathrm{d}\hat{x}$
$=$ $\frac{1}{2\pi\mathrm{i}}\int_{0}^{W/2}\mathrm{d}\hat{x}\,\chi(\hat{x},y)\int_{c-\mathrm{i}\infty}^{c+\mathrm{i}\infty}\mathrm{d}p\,e^{p(x-\hat{x})}$
$=$ $\frac{1}{2\pi\mathrm{i}}\int_{0}^{W/2}\mathrm{d}\hat{x}\,\chi(\hat{x},y)\int_{-\mathrm{i}\infty}^{\mathrm{i}\infty}\mathrm{i}\,\mathrm{d}\hat{p}\,e^{c(x-\hat{x})}\,e^{\mathrm{i}\hat{p}(x-\hat{x})}\quad(\text{putting}~{}p=c+\mathrm{i}\hat{p})$
$=$ $\frac{1}{2\pi}\int_{0}^{W/2}\mathrm{d}\hat{x}\,e^{c(x-\hat{x})}\chi(\hat{x},y)\int_{-\infty}^{\infty}\mathrm{d}\hat{p}\,e^{\mathrm{i}\hat{p}(x-\hat{x})},$
that can be rewritten in terms of the Dirac-delta function $\delta(x-\hat{x})=\frac{1}{2\pi}\int_{-\infty}^{+\infty}e^{\mathrm{i}\hat{p}(x-\hat{x})}\,\mathrm{d}\hat{p}$,
$\chi(x,y)=\frac{1}{2\pi}\int_{0}^{W/2}\mathrm{d}\hat{x}\,e^{c(x-\hat{x})}\chi(\hat{x},y)\,\delta(x-\hat{x}).$
We resolve the integral to the right-hand-side and we finally recover the Airy function in the domain of our sheet ($20\leq x\leq W/$ 2) and zero otherwise.
The Laplace transformation of the bi-harmonic equation (A4) is
$p^{4}\,f(y,p)+2\,p^{2}\,f_{,yy}(y,p)+f_{,yyyy}(y,p)=$ $p^{3}\,\chi|_{x=0}+2\,p^{2}\chi_{,x}|_{x=0}-p\left[-2\chi_{,yy}|_{x=0}+e^{-px}\chi_{,xx}|^{x=W/2}_{x=0}\right]$
$+2\chi_{,yyx}|_{x=0}-e^{-px}\chi_{,xxx}|^{x=W/2}_{x=0},$ (A13)
where we have already substituted the BCs at $x=W/2$ (A7b). Notice that we cannot infer all the boundary terms in the rhs of the equation solely from the Airy potential BCs (A7), and so we express the unknown functions with the aid of Fourier series. However, these terms should account for the stress singularities in the corners of a sheet when one edge is clamped and the other is free Williams (1952) (see Benthem (1963) and Xin and Davidovitch (2021a) for a complete calculation of the stress singularities in the vicinity of a sheet corner with clamped-free edges). For $\nu\sim 0.32$, $\sigma_{\theta\theta}$ and $\sigma_{\theta r}$ are given by Benthem (1963):
$\sigma_{\theta\theta}\left(\theta=\frac{\pi}{2}\right)=c_{1}r^{-1/4},\qquad\sigma_{r\theta}\left(\theta=\frac{\pi}{2}\right)=-c_{2}r^{-1/4},$ (A14)
@@
$\sigma_{xx}(x=0,y\to\pm 1)=c_{1}(1\pm x)^{-1/4},$ (A15a)
$\sigma_{xy}(x=0,y\to\pm 1)=\mp c_{2}(1\pm x)^{-1/4},$ (A15b)
where the choice for the signs follows from the following discussion. When stretching longitudinally the transversely clamped edge ($x=0$), we exert a tension onto the sheet. By the action-reaction principle, we expect that the sheet responds to this tension with a compressive stress, and hence the sign difference in the $\sigma_{xy}$ stress component. Along the transverse ($x$) direction, the transverse clamp at $x=0$ prevents the Poisson effect — by which a longitudinal stretching induces transverse contraction —, and hence it exerts a transverse compression to which the sheet responds generating a tensile transverse stress ($\sigma_{xx}(x=0,y\to\pm 1)>0$).
Hence, invoking the stress singularities at the corners (A15), we write
$\chi_{,yy}|_{x=0}=\sigma_{xx}(0,y)=c_{1}(1-x)^{-1/4}+c_{1}(1+x)^{-1/4}-c_{1}2^{-1/4}+\sum_{m}a_{m}\cos\left(\frac{m\pi}{2}y\right),$ (A16a)
$\chi_{,xy}|_{x=0}=-\sigma_{xy}(0,y)=c_{2}(1-x)^{-1/4}-c_{2}(1+x)^{-1/4}+c_{2}2^{-1/4}y+\sum_{n}b_{n}\cos\left(\frac{n\pi}{2}y\right),$ (A16b)
where $a_{m}$ and $b_{n}$ are yet to be determined constants, and $m=2k-1$, $n=2k$ for $k\in{\mathbb{Z}}^{+}$. We obtain from these the remaining unknown functions in (A13):
@@
$\chi_{,xyy}|_{x=0}=$ $-\sigma_{xy,y}(0,y)=\frac{c_{2}}{4}(1-x)^{-5/4}+\frac{c_{2}}{4}(1+x)^{-5/4}+c_{2}2^{-1/4}+\frac{\partial}{\partial y}\sum_{n}b_{n}\cos\left(\frac{n\pi}{2}y\right),$ (A17b)
@@
$\chi_{,x}|_{x=0}=$ $\int\mathrm{d}y\,\chi_{,xy}|_{x=0}+b_{0}=b_{0}-\frac{4}{3}c_{2}(1-x)^{3/4}-\frac{4}{3}c_{2}(1+x)^{3/4}+c_{2}2^{-5/4}x^{2}+\sum_{n}b_{n}\left(\frac{-2}{n\pi}\right)\cos\left(\frac{n\pi}{2}y\right),$ (A17d)
$\chi|_{x=0}=$ $\int\mathrm{d}y\,\left(\int\mathrm{d}\hat{y}\,\chi_{,\hat{y}}|_{x=0}+c_{4}\right)+c_{5}=c_{5}+c_{4}y+\frac{16}{21}c_{1}(1+x)^{7/4}-c_{1}2^{-5/4}x^{2}-\sum_{m}a_{m}\left(\frac{2}{m\pi}\right)^{2}\cos\left(\frac{m\pi}{2}y\right),$ (A17e)
where $c_{4}=0$, $c_{5}=c_{1}\left(-\frac{16}{21}2^{7/4}+2^{-5/4}\right)$ are computed using the BCs (A7a). Finally, we use the cosine Fourier series for $\chi_{,xx}|_{x=W/2}$ and $\chi_{,xxx}|_{x=W/2}$,
$\chi_{,xx}|_{x=W/2}=$ $\sum_{m}\alpha_{m}\cos\left(\frac{m\pi}{2}y\right),$ (A18a)
$\chi_{,xxx}|_{x=W/2}=$ $\sum_{m}\beta_{m}\cos\left(\frac{m\pi}{2}y\right),$ (A18b)
@@
-- $\nu c_{2}\left[\frac{1}{4}(1-y)^{-5/4}+\frac{1}{4}(1+x)^{-5/4}+2^{-1/4}\right]$
@@
-- $\nu\frac{\partial}{\partial y}\sum_{n}b_{n}\sin\left(\frac{n\pi}{2}y\right)-e^{-p\,W/2}\sum_{m}\beta_{m}\cos\left(\frac{m\pi}{2}y\right),$ (A19)
@@
We solve this differential equation for $f(y,p)$ and the corresponding BCs (A7a), which become
$f(\pm 1,p)=0;~{}f_{,y}(y,p)|_{y=\pm 1}=0,$ (A20)
we write the solution $f(y,p)$
@@
$+$ $\frac{2\,f_{p,y}(1,p)}{\sin 2p+2p}\left[\sin p\cos py-(\cos p)y\sin py\right],$ (A21)
with the particular integral
@@
-- $\sum_{n}b_{n}\frac{p^{2}\frac{2}{n\pi}+\nu\frac{n\pi}{2}}{\left[p^{2}-\left(\frac{n\pi}{2}\right)^{2}\right]^{2}}\cos\left(\frac{n\pi}{2}y\right),$ (A22)
$m=2k-1$, $n=2k$ for $k\in{\mathbb{Z}}^{+}$; where, to express this more compactly, we defined an ‘auxiliary’ function $R[q;p,y]$:
$R[q;p,\pm y]\equiv\frac{1}{2p^{3}}\int_{0}^{y}(1\pm\xi)^{q}\sin p(y-\xi)\,\mathrm{d}\xi-\frac{1}{2p^{2}}\int_{0}^{y}(1\pm\xi)^{q}(y-\xi)\cos p(y-\xi)\,\mathrm{d}\xi.$ (A23)
To obtain the Airy potential, $\chi(x,y)$, we compute the inverse Laplace transform (A12) of the function $f(y,p)$ (A21). To calculate the line integral (A12), we define a closed contour $\mathcal{C}$ in the complex plane and then apply the Residue Theorem Morse and Feshbach (1981). Note that, for arbitrary values of $(c_{1},b_{0},a_{m},b_{n},\alpha_{m},\beta_{m})$, the function $f(y,p)$ has poles at $p=p_{k}$, where $p_{k}$ are the (complex) roots of the trigonometric equation (A9) ($p_{k}\neq 0$) — the double poles at $p=\pm n\pi/2$, $p=\pm m\pi/2$, and the pole at $p=0$ can be removed. Thus, in order that the integral (A12) converges we need these poles at $p=p_{k}$ to be removed, i.e. their residues to vanish
$\mathrm{Res}[f(y,p),p=p_{k}]=\frac{1}{2}[f_{p,y}(y,p_{k})|_{y=1}-\sin^{2}p_{k}f_{p}(1,p_{k})]\frac{[\sin p_{k}\cos p_{k}y-(\cos p_{k})y\sin p_{k}y]}{\cos^{2}p_{k}}=0.$ (A24)
This leads to an infinite system of algebraic equations in the unknowns $(c_{1},b_{0},a_{m},b_{n},\alpha_{m},\beta_{m})$. Following the calculation of Benthem (1963), we take only a finite number of unknowns — i.e. we truncate the Fourier series in the particular integral (A22) —; the sufficient number of equations to calculate $(c_{1},b_{0},a_{m},b_{n},\alpha_{m},\beta_{m})$ is then obtained by using the necessary number of values of $p_{k}$ (ordered by increasing real part). Substituting the numerical values of $(c_{1},b_{0},a_{m},b_{n},\alpha_{m},\beta_{m})$ into the particular integral (A22), we obtain a function $f(y,p)$ that has no poles, and thus whose complex line integral (A12) converges.
So far our discussion has focused on the function $f(y,p)$, yet we notice that it multiplies the exponential $e^{px}$ in the inverse Laplace formula (A12). We close the integration contour so that $f(y,p)e^{px}$ is an analytic function, continuous within $\mathcal{C}$. Hence, by writing $f(y,p)$ as
$f(y,p)=e^{-pW/2}f_{+}(y,p)+f_{-}(y,p)$ , (A25a)
@@
$+$ $m\pi\sin\left(\frac{m\pi}{2}\right)\frac{\left[\sin p\cos py-(\cos p)y\sin py\right]}{\sin 2p+2p}\Bigg{]},$ (A25b)
$\text{and}\qquad f_{-}(y,p)=$ $f(y,p)-e^{-pW/2}f_{+}(y,p),$ (A25c)
$m=2k-1$ for $k\in{\mathbb{Z}}^{+}$, and substituting this into the integral (A12), we find that the integration line must be closed in either of the contours $\mathcal{C}_{\pm}$ in Fig. A9 for the integral of each term. Note that the functions $f_{\pm}(y,p)$ in (A25) have poles at $p=p_{k}$. Applying the Residue Theorem Morse and Feshbach (1981) (or Cauchy’s Theorem in those cases for which the integrand is $f(y,p)e^{px}$), we find that
$\chi(x,y)=\begin{cases}\frac{1}{2\pi\mathrm{i}}\oint\limits_{\mathcal{C}_{+}}f(y,p)e^{px}\,\mathrm{d}p=0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{for $x<0$},\\ \frac{1}{2\pi\mathrm{i}}\oint\limits_{\mathcal{C}_{+}}f_{+}(y,p)e^{p(x-W/2)}+\oint\limits_{\mathcal{C}_{-}}f_{-}(y,p)e^{px}\,\mathrm{d}p=\\ \qquad-\sum\limits_{k}e^{p_{k}(x-W/2)}\mathrm{Res}[f_{+}(y,p_{k}),\mathrm{Re}[p_{k}]>0]+\sum\limits_{k}e^{p_{k}x}\mathrm{Res}[f_{-}(y,p_{k}),\mathrm{Re}[p_{k}]<0],\quad\text{for $0\leq x\leq W/2$}\\ \frac{1}{2\pi\mathrm{i}}\oint\limits_{\mathcal{C}_{-}}f(y,p)e^{px}\,\mathrm{d}p=0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{for $x>W/2$}.\end{cases}$ (A26)
Note the different signs in the two terms in the Airy potential within the sheet domain ($x\in[0,W/2]$); "a closed contour is described in a positive direction with respect to the domain enclosed by the contour if with respect to some point inside the domain the contour is traversed in a counterclockwise directionMorse and Feshbach (1981).
@@
Extended Data Fig. A9: Contours in the complex plane a-b.

### v2

@@ Appendix E Details of Mathematical Modelling
We benchmark the expressions for the stress in the planar state (S10) with the finite-element simulations of a sheet ($L/W=2,~{}t/W=2.5\times 10^{-3},~{}W_{\mathrm{load}}/W=0.1,~{}E=3000~{}\text{MPa},~{}\nu=0.3$) subjected to longitudinal strain $\epsilon_{0}=0.025$. Considering the scaling assumed previously (i.e. lengths are rescaled by $L/2$ and $L=2$, and stresses are normalized by $T_{0}=Y\,\epsilon_{0}$), we plot in Fig. S9 the planar stress state obtained from (S10) and the numerical simulations (for which we consider only the region $x\geq W_{\mathrm{load}}/2$). We notice a good agreement between the predictions of the analytic model and the corresponding numerical simulations: compare Fig. S9a with Fig. S9b (transverse stress) and Fig. S9d with Fig. S9e (longitudinal stress). The transverse stress (Fig. S9a and b) is tensile at the corners where the characteristic tension is $T_{0}$, and compressive as we move away from the corners towards the centre — we have included the contour line at $\sigma_{xx}/T_{0}=0$ to evince the coexistence of compressive and tensile areas in the sheet. Close to the $y$-axis, the cross-sectional profiles of this stress component (Fig. S9c) help us visualize the tension-to-compression transition as we move away from the corners, and the stress singularity at their vicinity (see the next section for more details on the stress singularities at the corners of a sheet). Away from the $y$-axis, the transverse stress within the sheet fades rapidly. Conversely, the longitudinal stress (Fig. S9d and e) is tensile everywhere in the sheet, and hence is given less attention in our discussion.
@@
Extended Data Fig. S9: Stress in the planar state. a Analytic solution for the transverse stress (S10a) within the sheet subjected to a longitudinal localized load at the corners as per Fig. S8b; $W/L=0.45$ and we take $n=16$ terms in the series. b Numerical results for the transverse stress obtained by finite-element simulations of a sheet ($L/W=2,~{}t/W=0.0025,~{}W_{\mathrm{load}}/W=0.1,~{}E=3000~{}\text{MPa},~{}\nu=0.3$) subjected to a longitudinal strain $\epsilon_{0}=0.025$. The numerical results are rescaled according to the scaling in the mathematical model, and they are only shown within $x\geq W_{\mathrm{load}}/2$. The same colorbar as in a applies. c Cross-sectional profiles of the transverse stress at different values of $x$; symbols for the numerics, lines for the analytic expression. d Longitudinal component of the stress as predicted by (S10b). e Numerical results for the longitudinal stress with the same colorbar as in d.
@@
We further examine the transverse stress for different values of the aspect ratio $W/L$. We already noticed that the transverse stresses are weak away from the corners of the sheet that are loaded. Thus, we focus on the stress state close to the loaded corners. We plot in Fig. S12 the expression (S10a) shifted in $y$, $y\to y+1$, and we scale both axes by $W/2$. We observe a universal stress state for $W/L\lesssim 1/4$: the transverse stress is tensile at the vicinity of the corner, and compressive within an area that borders the narrow tensile part. The width of this compressing sector scales with $\sim W$ — we have included contour lines every $\Delta(\sigma_{xx}\/T_{0})=-0.05$ in the compressive part so as to evince the behavior of these stresses. This scaling is paramount in our analysis in the main text. For $W/L\gtrsim 1/4$, the compressive sectors corresponding to both corners overlap, and hence the sheet has a single compressed area spanning the centre of the sheet, while the two tensile sectors prevail at the corners.
@@
Extended Data Fig. S10: Transverse stress in the planar state for different aspect ratios. a-h Density plots of the stress by the expression (S10a). The longitudinal coordinate, $y$, is shifted $y\to y+1$; and both axes are scaled by $W/2$. The same colorbar (only shown in a) applies to all panels. The number of terms, $n$, taken in the series (S10a) is labeled in each graph. Contour lines are displayed at $\sigma_{xx}/T_{0}=(0,-0.05,-0.1,-0.15)$.
Benthem’s Laplace transform method
We define the function $f(y,p)$ as the Laplace transformation of the Airy potential,
$f(y,p)=\int_{0}^{W/2}\chi(x,y)\,e^{-px}\,\mathrm{d}x,$ (S11)
@@
$\chi(x,y)=\frac{1}{2\pi\mathrm{i}}\int_{c-\mathrm{i}\infty}^{c+\mathrm{i}\infty}f(y,p)\,e^{px}\,\mathrm{d}p\quad(c\in\mathbb{R},c>0).$ (S12)
The Laplace transformation of the bi-harmonic equation (S4) is
$p^{4}\,f(y,p)+2\,p^{2}\,f_{,yy}(y,p)+f_{,yyyy}(y,p)=$ $p^{3}\,\chi|_{x=0}+\,p^{2}\chi_{,x}|_{x=0}-p\left[-2\chi_{,yy}|_{x=0}+e^{-px}\chi_{,xx}|^{x=W/2}_{x=0}\right]$
$+2\chi_{,yyx}|_{x=0}-e^{-px}\chi_{,xxx}|^{x=W/2}_{x=0},$ (S13)
where we have already substituted the BCs at $x=W/2$ (S7b). Notice that we cannot infer all the boundary terms in the rhs of the equation solely from the Airy potential BCs (S7), and so we express the unknown functions with the aid of Fourier series. However, these terms should account for the stress singularities in the corners of a sheet when one edge is clamped and the other is free Williams (1952) (see Benthem (1963) and Xin and Davidovitch (2021a) for a complete calculation of the stress singularities in the vicinity of a sheet corner with clamped-free edges). For $\nu\sim 0.32$, $\sigma_{\theta\theta}$ and $\sigma_{\theta r}$ are given by Benthem (1963):
$\sigma_{\theta\theta}\left(\theta=\frac{\pi}{2}\right)=c_{1}r^{-1/4},\qquad\sigma_{r\theta}\left(\theta=\frac{\pi}{2}\right)=-c_{2}r^{-1/4},$ (S14)
@@
$\sigma_{xx}(x=0,y\to\pm 1)=c_{1}(1\mp y)^{-1/4},$ (S15a)
$\sigma_{xy}(x=0,y\to\pm 1)=\mp c_{2}(1\mp y)^{-1/4},$ (S15b)
where the choice for the signs follows from the following discussion. When stretching longitudinally the transversely clamped edge ($x=0$), we exert a tension onto the sheet, and hence $\sigma_{xy}(x=0,y\to\pm 1)$ should be pointing outwards at both corners: $\sigma_{xy}(x=0,y\to\pm 1)\to\mp\infty$ (notice that the surface’s unit vector carries an extra minus sign in the definition of the component of the stress tensor as it points downwards in the $x$-direction). Along the transverse ($x$) direction, the transverse clamp at $x=0$ prevents Poisson’s effect — by which a longitudinal stretching induces transverse contraction — and hence it exerts a tensile transverse stress, thus $\sigma_{xx}(x=0,y\to\pm 1)\to+\infty$.
Hence, invoking the stress singularities at the corners (S15), we write
$\chi_{,yy}|_{x=0}=\sigma_{xx}(0,y)=c_{1}(1-y)^{-1/4}+c_{1}(1+y)^{-1/4}-c_{1}2^{-1/4}+\sum_{m}a_{m}\cos\left(\frac{m\pi}{2}y\right),$ (S16a)
$\chi_{,xy}|_{x=0}=-\sigma_{xy}(0,y)=c_{2}(1-y)^{-1/4}-c_{2}(1+y)^{-1/4}+c_{2}2^{-1/4}y+\sum_{n}b_{n}\sin\left(\frac{n\pi}{2}y\right),$ (S16b)
where $a_{m}$ and $b_{n}$ are yet to be determined constants, and $m=2k-1$, $n=2k$ for $k\in{\mathbb{Z}}^{+}$. We obtain from these the remaining unknown functions in (S13):
@@
$\chi_{,xyy}|_{x=0}=$ $-\sigma_{xy,y}(0,y)=\frac{c_{2}}{4}(1-y)^{-5/4}+\frac{c_{2}}{4}(1+y)^{-5/4}+c_{2}2^{-1/4}+\frac{\partial}{\partial y}\sum_{n}b_{n}\sin\left(\frac{n\pi}{2}y\right),$ (S17b)
@@
$\chi_{,x}|_{x=0}=$ $\int\mathrm{d}y\,\chi_{,xy}|_{x=0}+b_{0}=b_{0}-\frac{4}{3}c_{2}(1-y)^{3/4}-\frac{4}{3}c_{2}(1+y)^{3/4}+c_{2}2^{-5/4}y^{2}+\sum_{n}b_{n}\left(\frac{-2}{n\pi}\right)\cos\left(\frac{n\pi}{2}y\right),$ (S17d)
$\chi|_{x=0}=$ $\int\mathrm{d}y\,\left(\int\mathrm{d}\hat{y}\,\chi_{,\hat{y}\hat{y}}|_{x=0}+c_{4}\right)+c_{5}$
$=$ $c_{5}+c_{4}y+\frac{16}{21}c_{1}(1-y)^{7/4}+\frac{16}{21}c_{1}(1+y)^{7/4}-c_{1}2^{-5/4}y^{2}-\sum_{m}a_{m}\left(\frac{2}{m\pi}\right)^{2}\cos\left(\frac{m\pi}{2}y\right),$ (S17e)
where $c_{4}=0$, $c_{5}=c_{1}\left(-\frac{16}{21}2^{7/4}+2^{-5/4}\right)$ are computed using the BCs (S7a). Finally, we use the cosine Fourier series for $\chi_{,xx}|_{x=W/2}$ and $\chi_{,xxx}|_{x=W/2}$,
$\chi_{,xx}|_{x=W/2}=$ $\sum_{m}\alpha_{m}\cos\left(\frac{m\pi}{2}y\right),$ (S18a)
$\chi_{,xxx}|_{x=W/2}=$ $\sum_{m}\beta_{m}\cos\left(\frac{m\pi}{2}y\right),$ (S18b)
@@
-- $\nu c_{2}\left[\frac{1}{4}(1-y)^{-5/4}+\frac{1}{4}(1+y)^{-5/4}+2^{-1/4}\right]$
@@
-- $\nu\frac{\partial}{\partial y}\sum_{n}b_{n}\sin\left(\frac{n\pi}{2}y\right)-e^{-p\,W/2}\sum_{m}\beta_{m}\cos\left(\frac{m\pi}{2}y\right),$ (S19)
@@
We solve this differential equation for $f(y,p)$ and the corresponding BCs (S7a), which become
$f(\pm 1,p)=0;~{}f_{,y}(y,p)|_{y=\pm 1}=0,$ (S20)
and we write the solution $f(y,p)$ as
@@
$+$ $\frac{2\,f_{p,y}(1,p)}{\sin 2p+2p}\left[\sin p\cos py-(\cos p)y\sin py\right].$ (S21)
Here the particular integral
@@
-- $\sum_{n}b_{n}\frac{p^{2}\frac{2}{n\pi}+\nu\frac{n\pi}{2}}{\left[p^{2}-\left(\frac{n\pi}{2}\right)^{2}\right]^{2}}\cos\left(\frac{n\pi}{2}y\right),$ (S22)
$m=2k-1$, $n=2k$ for $k\in{\mathbb{Z}}^{+}$; and, to express this more compactly, we defined an ‘auxiliary’ function $R[q;p,y]$:
$R[q;p,\pm y]\equiv\frac{1}{2p^{3}}\int_{0}^{y}(1\pm\xi)^{q}\sin p(y-\xi)\,\mathrm{d}\xi-\frac{1}{2p^{2}}\int_{0}^{y}(1\pm\xi)^{q}(y-\xi)\cos p(y-\xi)\,\mathrm{d}\xi.$ (S23)
To obtain the Airy potential, $\chi(x,y)$, we compute the inverse Laplace transform (S12) of the function $f(y,p)$ (S21). To calculate the line integral (S12), we define a closed contour $\mathcal{C}$ in the complex plane and then apply the Residue Theorem Morse and Feshbach (1981). Note that, for arbitrary values of $(c_{1},b_{0},a_{m},b_{n},\alpha_{m},\beta_{m})$, the function $f(y,p)$ has poles at $p=p_{k}$, where $p_{k}$ are the (complex) roots of the trigonometric equation (S9) ($p_{k}\neq 0$) — the double poles at $p=\pm n\pi/2$, $p=\pm m\pi/2$, and the pole at $p=0$ can be removed. Thus, in order that the integral (S12) converges we need these poles at $p=p_{k}$ to be removed, i.e. their residues to vanish
$\mathrm{Res}[f(y,p),p=p_{k}]=\frac{1}{2}[f_{p,y}(y,p_{k})|_{y=1}-\sin^{2}p_{k}f_{p}(1,p_{k})]\frac{[\sin p_{k}\cos p_{k}y-(\cos p_{k})y\sin p_{k}y]}{\cos^{2}p_{k}}=0.$ (S24)
This leads to an infinite system of algebraic equations in the unknowns $(c_{1},b_{0},a_{m},b_{n},\alpha_{m},\beta_{m})$. Following the calculation of Benthem Benthem (1963), we take only a finite number of unknowns — i.e. we truncate the Fourier series in the particular integral (S22); the number of equations sufficient to calculate $(c_{1},b_{0},a_{m},b_{n},\alpha_{m},\beta_{m})$ is then obtained by using the necessary number of values of $p_{k}$ (ordered by increasing real part). Substituting the numerical values of $(c_{1},b_{0},a_{m},b_{n},\alpha_{m},\beta_{m})$ into the particular integral (S22), we obtain a function $f(y,p)$ that has no poles, and thus whose complex line integral (S12) converges.
So far our discussion has focused on the function $f(y,p)$, yet we notice that it multiplies the exponential $e^{px}$ in the inverse Laplace formula (S12). We close the integration contour so that $f(y,p)e^{px}$ is an analytic function, continuous within $\mathcal{C}$. Hence, by writing $f(y,p)$ as
$f(y,p)=e^{-pW/2}f_{+}(y,p)+f_{-}(y,p)$ , (S25a)
@@
$+$ $m\pi\sin\left(\frac{m\pi}{2}\right)\frac{\left[\sin p\cos py-(\cos p)y\sin py\right]}{\sin 2p+2p}\Bigg{]},$ (S25b)
$\text{and}\qquad f_{-}(y,p)=$ $f(y,p)-e^{-pW/2}f_{+}(y,p),$ (S25c)
$m=2k-1$ for $k\in{\mathbb{Z}}^{+}$, and substituting this into the integral (S12), we find that the integration line must be closed in either of the contours $\mathcal{C}_{\pm}$ in Fig. S11 for the integral of each term. Note that the functions $f_{\pm}(y,p)$ in (S25) have poles at $p=p_{k}$. Applying the Residue Theorem Morse and Feshbach (1981) (or Cauchy’s Theorem in those cases for which the integrand is $f(y,p)e^{px}$), we find that
$\chi(x,y)=\begin{cases}\frac{1}{2\pi\mathrm{i}}\oint\limits_{\mathcal{C}_{+}}f(y,p)e^{px}\,\mathrm{d}p=0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{for $x<0$},\\ \frac{1}{2\pi\mathrm{i}}\oint\limits_{\mathcal{C}_{+}}f_{+}(y,p)e^{p(x-W/2)}+\oint\limits_{\mathcal{C}_{-}}f_{-}(y,p)e^{px}\,\mathrm{d}p=\\ \qquad-\sum\limits_{k}e^{p_{k}(x-W/2)}\mathrm{Res}[f_{+}(y,p_{k}),\mathrm{Re}[p_{k}]>0]+\sum\limits_{k}e^{p_{k}x}\mathrm{Res}[f_{-}(y,p_{k}),\mathrm{Re}[p_{k}]<0],\quad\text{for $0\leq x\leq W/2$}\\ \frac{1}{2\pi\mathrm{i}}\oint\limits_{\mathcal{C}_{-}}f(y,p)e^{px}\,\mathrm{d}p=0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{for $x>W/2$}.\end{cases}$ (S26)
Note the different signs in the two terms in the Airy potential within the sheet domain ($x\in[0,W/2]$); "a closed contour is described in a positive direction with respect to the domain enclosed by the contour if, with respect to some point inside the domain, the contour is traversed in a counterclockwise directionMorse and Feshbach (1981).
@@
Extended Data Fig. S11: Contours in the complex plane. The closed contours over which we apply the Residue Theorem in (S26) — these encircle all poles of the functions $f_{+}(y,p)$ (panel a) and $f_{-}(y,p)$ (panel b); the radii of circular portion of these contours must be chosen to be greater than the modulus of largest pole, $|p_{k}|$.